OpenStudy (anonymous):

6+x ___=4 2 how do you solve this?

OpenStudy (saifoo.khan):

\[\frac{6+x}{2}=4\] \[6+x=8\] \[x = 8 - 6\] x=2

OpenStudy (anonymous):

so you dont divide anything

OpenStudy (saifoo.khan):

No, we multiply two on both sides to get rid of the 2..

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

6+x=8 x=8-6 x=2

OpenStudy (anonymous):

THANKS THE BOTH OF YOU!!!

OpenStudy (anonymous):

OK SO 9+n __ 4 =15 do you multiply both sides by 4?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

not both side by 4

OpenStudy (anonymous):

just when 4 goes to 15 it multiply by it

OpenStudy (anonymous):

ok so how do i work this problem?

OpenStudy (anonymous):

9+n=15*4 9+n=60 n=60-9 n=51

OpenStudy (anonymous):

ok thanks so what about b+11 ___ 3=-2

OpenStudy (anonymous):

ary

OpenStudy (anonymous):

b+11=-2*3 b+11=-6 b=-17

OpenStudy (anonymous):

ok so for -12r+4=100 do you add 100+4 then divide by -12

OpenStudy (anonymous):

no

OpenStudy (anonymous):

subtrct

OpenStudy (anonymous):

oo because its a negitave

OpenStudy (anonymous):

100-4 96 then divid by -12

OpenStudy (anonymous):

no bc it positve when it goes to other hand it will negative

OpenStudy (anonymous):

why did you subtract

OpenStudy (anonymous):

reason is above

OpenStudy (anonymous):

yeah sorry you do inverse operation

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

what r u saying?

OpenStudy (anonymous):

im asking why did you subtract?

OpenStudy (anonymous):

because its positive when it goes to other hand it must be negative

OpenStudy (anonymous):

when any no goes to other hand its chang its sign

OpenStudy (anonymous):

if it positive then other hand its negative if negative then other hand it positive

OpenStudy (anonymous):

yeah so inverse operation

OpenStudy (anonymous):

if multiplt then divid on othr side

OpenStudy (anonymous):

if divid so multiply on other site

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

hve u got it now?